OpenStudy (anonymous):

the number of 3x3 non singular matrices with four entries as 1 and all other entries as 0 is

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (anonymous):

@dan815 ?

OpenStudy (anonymous):

@mathslover ?

OpenStudy (anonymous):

@vishweshshrimali5 ?

OpenStudy (anonymous):

@Miracrown ?

ganeshie8 (ganeshie8):

you cannot have a row or a column with all 0's

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

total ways of filling all elements =9!/4!5!

ganeshie8 (ganeshie8):

yes pick 4 elements from 9 : 9 choose 4

OpenStudy (anonymous):

ways in which 0 occupies rows and colums=3C1*2

ganeshie8 (ganeshie8):

may be we can try this : (9 choose 4) - ( total number of ways of having all 0's in a line)

OpenStudy (anonymous):

ways in which 0 occupies rows and colums=3C1*2?

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

for example : if u fill first row with "0 0 0" : there will be (6 choose 4) different combinations for selecting 1's on remaining two row elements

OpenStudy (anonymous):

there will be (6 choose 4) different combinations for selecting 1's on remaining two row elements??

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (anonymous):

?

OpenStudy (vishweshshrimali5):

@ganeshie8 Its not necessary that only matrices having only 0s in rows or columns will have zero determinant. Like 

ganeshie8 (ganeshie8):

Ahh true^ more cases to consider then !

ganeshie8 (ganeshie8):

this may work : "no two rows or columns should be same"

OpenStudy (vishweshshrimali5):

yeah

ganeshie8 (ganeshie8):

*add that constraint to previous

OpenStudy (vishweshshrimali5):

It would be better if we first find out the singular matrices and then subtract from all possible matrices.

ganeshie8 (ganeshie8):

1) no zeroes on a row/column 2) no two rows/columns should be same

ganeshie8 (ganeshie8):

yes im trying the same above^

OpenStudy (anonymous):

understood the constraints but how to find out them?

ganeshie8 (ganeshie8):

0 0 0 x x x x x x

ganeshie8 (ganeshie8):

you got four 1's to fill in those 6 x spots : (6 choose 4) ways

ganeshie8 (ganeshie8):

repeat for next two rows : 3*(6 choose 4) = 3(15) = 45 ways

ganeshie8 (ganeshie8):

similarly, working the columns give you another 45 ways but there will be few overlaps which you need to subtract

ganeshie8 (ganeshie8):

I think the overlap is 9; so the total number of ways a line(row/col) can have all zeroes is : 45 + 45 - 9 = 81 ways vishy check :)

OpenStudy (vishweshshrimali5):

One sec.

OpenStudy (vishweshshrimali5):

How did you obtain the overlap ?

ganeshie8 (ganeshie8):

0 0 0 1 1 0 1 1 0 thats a overlap : total permutations = 3 per each row => 9 in total

ganeshie8 (ganeshie8):

we got @mukushla here :)

OpenStudy (vishweshshrimali5):

Okay: 0 0 0 1 1 0 1 1 0 0 0 0 0 1 1 0 1 1 0 1 1 0 1 1 0 0 0 1 1 0 1 1 0 0 0 0 1 0 1 0 0 0 1 0 1 1 0 1 1 0 1 0 0 0 0 0 0 1 0 1 1 0 1 Any else I forgot ?

OpenStudy (vishweshshrimali5):

Okay 1 1 0 0 0 0 1 1 0 0 1 1 0 0 0 0 1 1

OpenStudy (vishweshshrimali5):

There we have 9

ganeshie8 (ganeshie8):

yes !

OpenStudy (vishweshshrimali5):

Yeah you are right till this step

ganeshie8 (ganeshie8):

good so we nailed one constraint, lets look at second constraint :)

ganeshie8 (ganeshie8):

@cody_123 are you wid us ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@vishweshshrimali5 ?

OpenStudy (vishweshshrimali5):

ohh sorry I was busy on another tab.

OpenStudy (vishweshshrimali5):

So did you understand the first constraint

OpenStudy (anonymous):

yes

OpenStudy (vishweshshrimali5):

Sorry cody I would be unable to respond you for atleast half and hour. I have some serious work to attend.

OpenStudy (vishweshshrimali5):

Sorry

OpenStudy (anonymous):

some hint

OpenStudy (anonymous):

@ParthKohli ?

OpenStudy (anonymous):

@dan815 ?

OpenStudy (anonymous):

@Miracrown

OpenStudy (anonymous):

if we have 2 columns same number of matrices=3*3=9

OpenStudy (anonymous):

@ganeshie8 ?

ganeshie8 (ganeshie8):

hey

OpenStudy (anonymous):

if we have 2 columns same number of matrices=3*3=9?

OpenStudy (anonymous):

how did y get the columns?

OpenStudy (anonymous):

if we fix one of the column with only 1's then there are 9 matrices

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

w8 a second how can u arrange all 1's in one column that would violate the condition of 4 1's

ganeshie8 (ganeshie8):

we made a mistake - lets work this thing again :)

OpenStudy (anonymous):

0 0 1 0 0 1 = 3 matrices 1 1 0 0 0 1 1 1 0 = 3 matrices 0 0 1 1 1 0 0 0 1 =3 matrices 0 0 1

ganeshie8 (ganeshie8):

Wow ! looks perfect !!

ganeshie8 (ganeshie8):

will there be any more matrices with same rows ?

OpenStudy (anonymous):

no all matrices will be over counted then

ganeshie8 (ganeshie8):

i think you're right ! do u have answer for this ? :)

OpenStudy (anonymous):

answer should be 126-(81+9)=36?

ganeshie8 (ganeshie8):

I think so it should be \(\le 36\) Need to confirm we did not miss counting any

OpenStudy (anonymous):

please see i dont think so there is any

OpenStudy (anonymous):

@ganeshie8 ?

ganeshie8 (ganeshie8):

@dan815 got some time ?

OpenStudy (anonymous):

@hartnn ?

OpenStudy (mathmate):

I think we have to fill the diagonal first, if we don't want a singular matrix.

OpenStudy (anonymous):

i think those matrices are counted

OpenStudy (mathmate):

So there is only one "1" left for the other 6 places.

OpenStudy (mathmate):

@cody_123 What do you mean by "counted"? :)

OpenStudy (anonymous):

i mean matrices where all the diagonal elements are same are counted in the above cases like 0 1 1 0 0 1 1 1 0

OpenStudy (mathmate):

If they are identical, they should be counted as one. But... When the matrix has a zero on the diagonal, it will be singular. Since the matrix has to be non-singular, all the diagonal terms must be non-zero. This will therefore take care of three "1"s. You can then place the remaining "1" wherever you want.

OpenStudy (anonymous):

so it is another constraint?

OpenStudy (mathmate):

Yes, it is a constraint that helps you. Because you know where three of the '1"s go, and leaves you with only one left in the other six places.